INAF queue observing, 2020 Jan 30 - Jan 31

Observers: A. Rossi, S. Paiano & Telescope Operator: S. Allanson & Support Astronomer: D. Thompson & Instrument: MODS1-2

At sunset the sky is partly cloudy but it is clearing up There is a little bit of moon, so the sky is brigther.

MODS: both ok

LBC:
  • Thanks to the great work of LBTO now LBC blue is available again, although is just following the guiding of LBCred.

LUCI:
  • LUCI1: not available all week
  • LUCI2: not mounted due to AGW upgrade for SOUL

Temperature = -6 Humidity = 30 Wind =

Summary

Night duration (12 degree - twilight) 1:30 - 13:30 UT = 12h 00m = 43200 seconds
Observing time: h min (97%)

Weather time loss 20min (3%) sec

Technical time loss: 0min (0%) sec

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Antoniucci V118Ori MODS1-2 dual, slit 0.8" 40.6/0.3h = mods1b.03-06 ,mods1r.09-12 mods2b.03-06, mods2r.06-09 = passing clouds completed
Antoniucci AGLL322 MODS1-2 dual, slit 0.8" 40.6/0.3h = mods1b.0007-0010 ,mods1r.0017-0020 mods2b.0007-0010, mods2r.0014-0017 = passing clouds completed
Antoniucci 2MASSJ06 MODS1-2 dual, slit 0.8" 40.6/0.3h = mods1b.0011-0014 ,mods1r.002528- mods2b.0011-14, mods2r.0021-0024 = passing clouds completed
Giunta/DDT_BACKUP_1 fieldG LBCr,LBCb g,r 2h   seeing variable 1.3-2.5
Mannucci J1045 MODS1-2 dual, 0.6"     3 times
Mannucci J1115 MODS1-2 dual, 0.6"     1 time
Mannucci J1306 MODS1-2 dual, 0.6"     3 times
Calibration flat MODS1-2 red slitless     per Belladitta

Details

01:58 We start with the Antoniucci's program pointing V1118Ori with MODS1-2.
  • Seeing = 1.2"
  • Acq : = mods1r.20200131.0003-0008 mods2r.0003-0005 = Difficult to find the source in the slit during the acquisition because to high cloudiness (>2 mag of exctintion)
  • Sci : = mods1b.03-06 ,mods1r.09-12 mods2b.03-06, mods2r.06-09 =
  • 02:35 clouds are gone, seeing 1.5" on DIMM
  • 02:50 seeing 1.0" on DIMM
  • 02:55 passing cloud, 2 mag extinction, just for 5 min

03:09 Slew to Antoniucci AGLL322
  • Acq : = mods1r.0013-0016 mods2r.0010-0013 =
  • Sci : = mods1b.0007-0010 ,mods1r.0017-0020 mods2b.0007-0010, mods2r.0014-0017 =
  • 03:25 DIMM seeing 1.5" variable
  • 03:45 DIMM seeing 2"

*03:55 Slew to Antoniucci 2MASSJ06
  • Seeing : 1.45
  • Acq : = mods1r.0021-0022 mods2r.0018-0019 =
  • Sci : = mods1b.0011-0014 ,mods1r.002528- mods2b.0011-14, mods2r.0021-0024 =
  • 04:28 DIMM seeing 1.2" variable4

04:38 Switch to LBC
  • DIMM seeing >1.3"

05:05 Slew to Alessio program DDT_BACKUP_1
  • dohybrid
  • seeing >2, difficult to converge. we try dofpia, X2
  • copointing
  • send DDT_BACKUP_1_1_gilli_field1.xml
  • seeing >2 on images and ~1.5" on DIMM
  • 05:46 send DDT_BACKUP_1_3_gilli_field2.xml
  • seeing ~2
  • 06:00 dopia, X2 " il red non ha trovato una appropriata collimazione"
  • copointing
  • 06:21 send DDT_BACKUP_1_5_gilli_field3.xml
  • stars elongated
  • dopia again
  • 06:38 DDT_BACKUP_1_7_gilli_field4.xml
  • seeing 1.5" on images
  • 06:47 DDT_BACKUP_1_9_gilli_field5.xml
  • 07:00 DDT_BACKUP_1_11_gilli_field6.xml
  • 07:10 DDT_BACKUP_1_13_gilli_field7.xml
  • seeing 1.4"
  • dofpia
  • 07:26 DDT_BACKUP_1_15_gilli_field8.xml
  • 07:35 DDT_BACKUP_1_17_gilli_field9.xml
  • we do again the first fields because seeing was >2
  • 07:44 DDT_BACKUP_1_1_gilli_field1.xml again
  • copoining
  • dofpia
  • 08:05 send DDT_BACKUP_1_3_gilli_field2.xml
  • seeing1.5"-2"
  • 08:15 send DDT_BACKUP_1_5_gilli_field3.xml

08:25 The seeing is still very bad, so we decide to switch to MODS

08:55 Slew to Mannucci J1045 PA 270
  • seeing 2" on DIMM
  • Acq : mods1r.0029-0032, mods2r.0025-0027
  • 09:15 Sci : mods1b.0015-16 ,mods1r.0033-34, mods2b.0015-16, mods2r.0028-29
  • we take 1 exp more becuse seing ~2 or more

09:35 Slew to Mannucci J111501 PA -10
  • seeing 2" on DIMM
  • Acq : mods1r.0035-0038 mods2r.0030-0033
  • seeing ~3" !!
  • 09:57 Sci : mods1b.0017-18 ,mods1r.0039-40, mods2b.0017-18, mods2r.0034-35

10:13 Slew to Mannucci J130646 PA 220
  • seeing 1.7" on DIMM
  • Acq : mods1r.0041-44 mods2r.0036-0038
  • Sci : mods1b.0019-20 ,mods1r.0045-46, mods2b.0019-20, mods2r.0039-40

10:48 Slew to Mannucci J1045 PA 270
  • seeing 2" on DIMM
  • Acq : mods1r.0047-49 mods2r.0041-0043
  • Sci : mods1b.0021-22 ,mods1r.0050-51, mods2b.0021-22, mods2r.0044-45

11:23 Slew to Mannucci J111501 PA +30
  • seeing >3" on DIMM* seeing too high we have to abort

11:30 Slew to Mannucci J130646 PA 220 (higher field probably better)
  • Acq : mods1r.0053-0055 mods2r.0047-0049
  • Sci : mods1b.0023 ,mods1r.0056, mods2b.0023, mods2r.0050
  • seeing 1.7" on DIMM

12:00 Slew to Mannucci J1045 PA 270
  • Acq : mods1r.0057-0059 mods2r.0051-0053
  • seeing 2-3" on DIMM
  • Sci : mods1b.0024-25 ,mods1r.0060-61, mods2b.0024-25, mods2r.0054-55

12:34 Slew to Mannucci J130646 PA 220
  • Acq : mods1r.0062-64 mods2r.0056-58
  • Sci : mods1b.0026-27 ,mods1r.0065-66, mods2b.0026-27, mods2r.0059-60
  • seeing 2-3" on DIMM

13:10 Slew to Mannucci J1045 PA 270
  • seeing too high cannot guide aborted, TO try to solve then we try again
  • Acq : mods1r.0068-70- mods2r.0062-64
  • we do not see anything in the aacquisition we stop

MODS CALIBRATION

FINE
Topic revision: r4 - 02 Feb 2020, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback